OpenStudy (anonymous):

log base 2 fifth root of 32 ?!?!?!?!?!??!?!?!?!!?!??! PLEASE HELP

OpenStudy (xishem):

\[\log_{2} \sqrt[5]{32}=\log_{2}2=1\]

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

how did you get log base 2 then 2?

OpenStudy (anonymous):

the fifth root of 32 is 2, because \[2^5=32\]

OpenStudy (anonymous):

then shouldn't the answer be five ?

OpenStudy (anonymous):

so your job is to find \[\log_2(2)\] which asks the question, "what power do you raise the number "2" to to get an answer of 2?" and the answer is clearly 1

OpenStudy (anonymous):

\[\log_2(2^5)=5\] yes

OpenStudy (anonymous):

but you have \[\log_2(2)\]

OpenStudy (anonymous):

wait..what?!

OpenStudy (anonymous):

\[2^{y} = 2^{5} \]

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

if your question was \[\log_2(2^5)\] the answer would certainly be 5 correct?

OpenStudy (anonymous):

i.e. \[\log_2(32)=5\]

OpenStudy (anonymous):

wait one sec let me soak this in

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so you're telling me that 2^5 does not equal fifth root of 32?

OpenStudy (anonymous):

yes i am. 2^5 = 32 not the fifth root of 32

OpenStudy (anonymous):

so what is it? and how would u find what it is ?

OpenStudy (anonymous):

i want to know the fifth rote of 32

OpenStudy (anonymous):

but since 2^5 = 32 then we know that \[\sqrt[5]{32}=2\]

OpenStudy (anonymous):

it is 32^ 1/5

OpenStudy (anonymous):

you can think of it that way, yes

OpenStudy (anonymous):

im not understanding what you're telling me

OpenStudy (anonymous):

which means \[(32)^{\frac{1}{5}}=(2^5)^{\frac{1}{5}}=2^1=2\]

OpenStudy (anonymous):

in english: two the the fifth power is 32, therefore the fifth root of 32 is 2

OpenStudy (anonymous):

i think i understand that. thank you very much. can you help me with a few more?

OpenStudy (anonymous):

post i will reply

OpenStudy (anonymous):

\[3 \log _{7} \sqrt[6]{49}\]

OpenStudy (anonymous):

ok lets do this the easy way. first of all \[\log(x^n)=n\log(x)\]

OpenStudy (anonymous):

yep. i understand tha : - )

OpenStudy (anonymous):

and we can rewrite \[\sqrt[6]{49}\] as \[49^{\frac{1}{6}}\] yes?

OpenStudy (anonymous):

ummm im kind of confused on those

OpenStudy (anonymous):

is that just a general rule ?

OpenStudy (anonymous):

you need to understand exponential notation for logs, because logs are exponents.

OpenStudy (anonymous):

yes for example \[\sqrt{x}=x^{\frac{1}{2}}\] \[\sqrt[3]{x}=x^{\frac{1}{3}}\] etc

OpenStudy (anonymous):

and in general \[\sqrt[n]{x}=x^{\frac{1}{n}}\]

OpenStudy (anonymous):

okay awesome. just wrote that on my rules sheet XD

OpenStudy (anonymous):

so we start with \[3\log_7(49^{\frac{1}{6}})\] and then bring the 1/6 right out front as a multiplier (coefficient)

OpenStudy (anonymous):

\[\frac{1}{6}\times 3\log_7(49)\] \[\frac{1}{2}\log_7(49)\] and idea what to do next?

OpenStudy (anonymous):

i really have no idea :(

OpenStudy (anonymous):

how can you write 49 with an exponent?

OpenStudy (anonymous):

what would u do if you have a number in front of the log? like the 1/2

OpenStudy (anonymous):

hold on, lets just take care of the 49 first. we can do this problem a different way second

OpenStudy (anonymous):

we would write \[49=7^2\]

OpenStudy (anonymous):

and then you have \[\frac{1}{2}\log_77^2)\] which should be easy now

OpenStudy (anonymous):

how would we incorporate the 1/2?

OpenStudy (anonymous):

since \[\log_b(b^n)=n\]

OpenStudy (anonymous):

but its 1/2 log

OpenStudy (anonymous):

we leave it for a moment. the one half out front is just a number, let it sit there

OpenStudy (anonymous):

so the answer is 2

OpenStudy (anonymous):

we get \[\frac{1}{2}\log_7(7^2)=\frac{1}{2}\times 2\]

OpenStudy (anonymous):

the answer to \[\log_7(7^2)=2\] and then \[\frac{1}{2}\times 2=1\]

OpenStudy (anonymous):

oh okay!!!!! wow i get that!

OpenStudy (anonymous):

if you like we can do this problem a different way, but maybe one way is enough, you tell me

OpenStudy (anonymous):

answer is 1

OpenStudy (anonymous):

yes it is 1

OpenStudy (anonymous):

i like this way. do u mind helping me with more?

OpenStudy (anonymous):

go ahead

OpenStudy (anonymous):

\[50 \log 5 \sqrt{125} \]

OpenStudy (anonymous):

so we would have

OpenStudy (anonymous):

5 ^ x = square route 125 ^ 50?

OpenStudy (anonymous):

oh yikes leave the 50 out front!

OpenStudy (anonymous):

i mean you are right, but leave it there to make your life easier. think of how to write 125 as a number raised to a power

OpenStudy (anonymous):

okay. lol so i have 50 log base 5^x = 5^ 3

OpenStudy (anonymous):

i believe that is wrong

OpenStudy (anonymous):

close, you just forget the square root

OpenStudy (anonymous):

but i don't have the square route anymore i got it al out of the square rote

OpenStudy (anonymous):

\[125=5^3\] and so \[\sqrt{125}=5^{\frac{3}{2}}\]

OpenStudy (anonymous):

i dont understand that

OpenStudy (anonymous):

just a moment ago we found out that \[\sqrt{x}=x^{\frac{1}{2}}\] right?

OpenStudy (anonymous):

\[\sqrt{125}=125^{\frac{1}{2}}=(5^3)^{\frac{1}{2}}=5^{\frac{3}{2}}\]

OpenStudy (anonymous):

yes. oh okay i understand

OpenStudy (anonymous):

so then what would we have

OpenStudy (anonymous):

so now we have \[50\log_5(5^{\frac{3}{2}})\] \[50\times \frac{3}{2}\] etc

OpenStudy (anonymous):

as soon as you say \[\log_b(\text{whatever})\] you should think of ways to write "whatever" as a power of b

OpenStudy (anonymous):

*see

OpenStudy (anonymous):

oh okay!!! wow!

OpenStudy (anonymous):

can u give me an example to try?

OpenStudy (anonymous):

\[\log_3(\sqrt[5]{27})\]

OpenStudy (anonymous):

can u do one with a number in front of the log

OpenStudy (anonymous):

the number out front does not really make any difference, but if you like i can write \[10\log_3(\sqrt[5]{27})\]

OpenStudy (anonymous):

okay let me get my paper lol

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

OMG

OpenStudy (anonymous):

really? i got it right?

OpenStudy (anonymous):

don't be so shocked, you will get the hang of it

OpenStudy (anonymous):

can u help me with more? lol

OpenStudy (anonymous):

after a couple practice problems you will think "not much to this"

OpenStudy (anonymous):

ok we can do another before i go

OpenStudy (anonymous):

\[\log _{3} 4 - 1/2 \log _{3} (6x-5)\]

OpenStudy (anonymous):

holy moly what are the instruction, write as a single log?

OpenStudy (anonymous):

lol i have no idea what to do at all

OpenStudy (anonymous):

first you need to make sure what the instructions are. what does it say before the problem?

OpenStudy (anonymous):

all it says is : solve

OpenStudy (anonymous):

maybe u an help me with a different one instead?

OpenStudy (anonymous):

you cannot "solve' because there is not an equation here. maybe it says "write as one log"

OpenStudy (anonymous):

which we can do if you like